A fair spinner has 9 equal sections: 3 red, 2 blue and 4 green.

It is spun twice.

What is the probability of getting blue both times?


1
Expert's answer
2021-10-04T18:27:58-0400

The probability of a blue section appearing every time is

"p = \frac{2}{{3 + 2 + 4}} = \frac{2}{9}"

Then the wanted probability is

"P = {p^2} = {\left( {\frac{2}{9}} \right)^2} = \frac{4}{{81}}"

Answer: "P = \frac{4}{{81}}"